what is the volume of 15.8 M nitric acid is needed to ensure 1.00 g sample copper completely reacts


1
Expert's answer
2021-10-04T05:09:59-0400

M(Cu) = 63.54 g/mol

n(Cu) "= \frac{1.00}{63.54} = 0.0157 \;mol"

According to the reaction equation:

n(HNO3) = 4n(Cu) "= 4 \times 0.0157 = 0.0629 \;mol"

Proportion:

15.8 mol – 1000 mL

0.0629 mol – x

"x = \frac{0.0629 \times 1000}{15.8} = 3.98 \; mL ≈ 4 \;mL"

Answer: 4 mL